OpenStudy (anonymous):

The formula for degrees Celsius is: C = 5 over 9 (F - 32), where F stands for degrees Fahrenheit. Part 1: Solve the formula for F. Show all steps. Part 2: Determine how many degrees Fahrenheit 2 degrees Celsius is.

OpenStudy (anonymous):

Part 1: Solve the formula for F. Show all steps. C = 5/9 (F – 32) 9C = 5(F - 32) F - 32 = 9C/5 F = 9C/5 + 32 -------------- Part 2: Determine how many degrees Fahrenheit 6 degrees Celsius is. F = 9*6/5 + 32 F = 54/5 + 32 F = 10.8 + 32 F = 42.8 ----------------------- To convert "in your head": Add 40 If converting from C to F, multiply by 1.8 If converting from F to C, divide by 1.8 Subtract 40 ------------- eg from 25C to F Add 40 --> 65 Multiply by 9/5 --? 585/5 = 117 Subtract 40 --> 77F

OpenStudy (anonymous):

did u understand?

OpenStudy (anonymous):

this time

OpenStudy (anonymous):

Yes thank you! :)

OpenStudy (anonymous):

thanx 4 the medal

OpenStudy (anonymous):

Your Welcome! :)

OpenStudy (anonymous):

do u have anymore questions?

OpenStudy (anonymous):

Thx for mine too. :)

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

Nope thanks for asking :)

OpenStudy (anonymous):

oh ok